OpenStudy (anonymous):

what is 10x+1+9x-11=180

OpenStudy (luigi0210):

it is an unsolved equation

OpenStudy (goldphenoix):

\[10x+1+9x-11=180\] Simplify: 19x -10 = 180 +10 +10 ----------------- 19x = 190 Divide: x is equal to 10. Hope this help!

OpenStudy (anonymous):

yes gold

OpenStudy (anonymous):

thank you

OpenStudy (goldphenoix):

No problem. I hope you have a nice day/afternoon/night. :)

OpenStudy (anonymous):

you tooooooooo